# Identifying Quantum Mechanical Statistics in Italian Corpora

### v1

@@ 5 Quantum statistics and meaning
The way in which meaning is attributed in the combination, or composition, of two concepts, i.e. two conceptual entities in our quantum theoretical framework (see Sect. 2), closely resembles the formation of entanglement in the composition, or combination, of two physical entities. In both cases, indeed, a phenomenon of ‘contextual updating’ occurs, which we have recently formulated (Aerts et al.,, 2024, 2023, 2023). Thus, let us illustrate what we mean by contextual updating and firstly consider the combination of two words, and the two corresponding concepts, in a piece of text produced by human language.
@@
Hence, the phenomenon of contextual updating can explain the way in which meaning attaches to and intertwines with the words in the composition of a piece of text. Moreover, in the process of writing, the writer has typically the intention of making the piece of text as clear as possible. In our quantum theoretical framework, this corresponds to the preparation of a pure entangled state, whose von Neumann entropy is equal to zero. Thus, in the writing up process, the words, thus the conceptual entities corresponding to them, always entangle together, namely, ‘collaborate’ with the aim of reducing the von Neumann entropy, hence, the overall uncertainty of the text, and producing globally a pure state (Aerts et al.,, 2023, 2023). More specifically, in this collaborative dynamics to produce entanglement, identical words, i.e. cognitons in the same micro-states, tend to clump, or cluster, together, thus breaking the statistical independence. It is this dynamics of meaning formation that would be responsible of the appearance of the quantum Bose–Einstein statistics, in human language.

### v2

@@ 5 Quantum statistics and meaning
The way in which meaning is attributed in the combination, or composition, of two concepts, i.e. two conceptual entities in our quantum theoretical framework (see Sect. 2), closely resembles the formation of entanglement in the composition, or combination, of two physical entities. In both cases, indeed, a phenomenon of ‘contextual updating’ occurs, which we have recently formulated (Aerts et al.,, 2024; Aerts et al., 2023a, ; Aerts et al., 2023b, ). Thus, let us illustrate what we mean by contextual updating and firstly consider the combination of two words, and the two corresponding concepts, in a piece of text produced by human language.
@@
Hence, the phenomenon of contextual updating can explain the way in which meaning attaches to and intertwines with the words in the composition of a piece of text. Moreover, in the process of writing, the writer has typically the intention of making the piece of text as clear as possible. In our quantum theoretical framework, this corresponds to the preparation of a pure entangled state, whose von Neumann entropy is equal to zero. Thus, in the writing up process, the words, thus the conceptual entities corresponding to them, always entangle together, namely, ‘collaborate’ with the aim of reducing the von Neumann entropy, hence, the overall uncertainty of the text, and producing globally a pure state (Aerts et al., 2023a, ; Aerts et al., 2023b, ). More specifically, in this collaborative dynamics to produce entanglement, identical words, i.e. cognitons in the same micro-states, tend to clump, or cluster, together, thus breaking the statistical independence. It is this dynamics of meaning formation that would be responsible of the appearance of the quantum Bose–Einstein statistics, in human language.